OpenStudy (anonymous):

HELP!!!!!!! multiply x^2/5 * x^2/9 the possible answers are x to the 2 sevenths power x to the 28 over 45 power x to the 8 over 45 power x to the 4 over 45

OpenStudy (anonymous):

\[\frac{ x^2 }{ 5 }*\frac{ x^2 }{ 9 } = \frac{ x^4 }{ 45 } \] Is that what you mean?

OpenStudy (anonymous):

no like x^2/5 times x^2/9

OpenStudy (anonymous):

howd you get 28?

OpenStudy (anonymous):

no one minute.. gotta redo \[x^(2/5) * x^(2/9) = x^(28/45)\]

OpenStudy (anonymous):

thats like 2/5 + 2/9 = 18/45 + 10/45 now thats right

OpenStudy (anonymous):

okay what about Which of the following represents 6 x to the 3 fourths power in radical form? cubed root of 6 x to the fourth power fourth root of 6 x cubed 6 cubed root of x to the fourth power 6 fourth root of cubed

OpenStudy (anonymous):

\[x^a*x^b = x^(a+b)\]

OpenStudy (anonymous):

Which of the following represents 6 x to the 3 fourths power in radical form? cubed root of 6 x to the fourth power fourth root of 6 x cubed 6 cubed root of x to the fourth power 6 fourth root of cubed

OpenStudy (anonymous):

\[(6x)^{3/4} = \sqrt[4]{(6x)^{3}} \] Is that what you mean?

OpenStudy (anonymous):

\[x ^{a/b} = \sqrt[b]{x ^{a}}\]